OpenStudy (anonymous):

25 squared divided by 5 cubed how would i write this out? I know it equals 5 but how would i write it??

OpenStudy (anonymous):

I was just about to say lmao it is 5. lol I just dont know how to write it out lol

OpenStudy (anonymous):

\[\frac{25^2}{5^3}\]

OpenStudy (anonymous):

or \[25^2 \div 5^3\]

OpenStudy (anonymous):

I got it right then lol

OpenStudy (jhannybean):

\[\large \frac{25^2}{5^3}\]\[\large \frac{(5^2)^2}{5^3}= \frac{5^4}{5^3}= 5^{4-3}= ?\]

OpenStudy (jhannybean):

that's another way to solve it, since 25^2 = 625 and 5^4 = 625.